### v1

@@ 3 Modeling methods
The labeled dataset and its corresponding image and weather data are used to predict road surface conditions across NYS. Classification ML algorithms, including CNNs and RFs, are trained to categorize each observation into one of five road surface classes for the Surface Condition Model. For the Obstruction Detection Model, a CNN is utilized to classify each observation as either obstructed or non-obstructed. Lastly, ensembling methods are employed to utilize the predictions from multiple models trained on different datasets, as discussed in Section 2.2. For all ML methods in this section, input data are subject to relevant normalization or preprocessing prior to training. Balanced class weighting is used to address the imbalanced dataset (Table 1), which is important to encourage the model to learn all classes rather than simply learning to predict majority classes, like wet and dry.
@@ 3.1 Image classification
Images contain a wide range of colors, textures, edges, and nonlinear, complex patterns. CNNs are commonly used image models because they can automatically extract these complex features in images to learn a classification task. Spatial relationships that exist in the images are learned through filters and convolutional layers (Chase et al., 2023), and nonlinearity in the data is learned by filters and convolutional layers with activation functions. Many different CNN models are trained to identify the best performing architectures and hyperparameters for this task.
The CNN models are trained on NVIDIA A100 and V100 GPUs using the Tensorflow python package (TensorFlow Developers, 2024). Images are cropped to remove the top 20% of the image that often contains text (e.g., timestamps or location, examples in Table 1) or is not focused on roadways (e.g., sky or distant scenery). Images are resized to 224x224 and encoded as RGB. Cropping, resizing, and standardization of the data are preprocessing steps that are done on the fly during data loading. The model weights from the epoch achieving highest validation accuracy are saved as the product of training. Table 2 shows additional CNN modeling details.
@@ 3.2.1 CNN models and hyperparameter tuning
Six CNN architectures, which range in complexity, include DenseNet121 from Huang et al. (2016), InceptionV3 from Szegedy et al. (2014), Mobilenet from Howard et al. (2017), ResNet50 from He et al. (2015), VGG16 from Simonyan and Zisserman (2015), and Xception from Chollet (2017). Data augmentation is used to diversify the training set by applying transformations to the original images. Such transformations include adjusting the brightness and/or contrast in the image, and horizontally flipping the image. These augmentations are made automatically and randomly to images on the fly during model training, with replacement, to make the dataset more robust and varied (TensorFlow Developers, 2024).
@@
Transfer learning involves using pretrained models where the backbone of the architecture (which include convolutional layers) is frozen; the weights are already learned from the commonly used ImageNet dataset (Deng et al., 2009). The Keras library provides both the architectures and pre-trained weights that are used for this work (Chollet and et al., 2015). The final layers of the architecture are the dense layers that ultimately result in the final classification, and will be referred to as the “top” of the model. The top is always trained specifically on the NYSDOT dataset. Two different architecture tops are considered: 1) the generic top (referred to as “GEN”) consists of a 2D global average pooling layer followed by three dense layers of size 1024, 512, and six (the output layer), and 2) the architecture-specific top (referred to as “AST”) consists of layers that align with the original architecture. The third consideration is the exclusion of transfer learning meaning that all weights from all model layers are trained using the NYSDOT dataset. Thus, the use of transfer learning with two different tops (GEN or AST), as well as the exclusion of any transfer learning, comprise three choices related to transfer learning that are considered in the baseline search.
@@
For hyperparameter tuning, L2 regularization and dropout are common regularization techniques used to reduce overfitting in CNN models (Géron, 2020). These regularization techniques are used on the top dense layers, and hyperparameter tuning is completed to identify the best performing L2 regularization weight and dropout rate. For the baseline models, these hyperparameters are set to 0. The hyperparameter space, shown in Table 3, includes four different L2 regularization weights ranging from 0.2 to 0.8, and four dropout rates ranging from 1E-7 to 0.1, resulting in 16 sets. Thus, including the baseline runs, there are 17 hyperparameter sets tested in total.
@@
The modeling process begins with running the 36 baseline models, from which the top four are selected using the model selection criteria discussed in Section 3.2.3. The four models then undergo hyperparameter tuning, and one top performing hyperparameter set is selected for each. Thus, four final CNN models for image classification are considered and carried over into the next step of the modeling process, the downstream classification ML model, which incorporates information from the CNN as well as weather data, as discussed in the next section. The final CNN is selected following hyperparameter tuning of the downstream model.
@@
Table 3: CNN model considerations. Baseline models are run (36 combinations), and then the top baseline models undergo hyperparameter tuning. The red text indicates the final best CNN model that is chosen and discussed in Section 4.1.
@@
Architecture Transfer Learning Augment. Dropout L2
Densenet, Inception, Mobilenet, Resnet, VGG16, Xception None         Transfer Learn: AST Transfer Learn: GEN Aug.      No Aug. 0.0, 0.2, 0.4, 0.6, 0.8 0.0, 1E-7, 1E-5, 0.001, 0.1
@@ 3.2.2 Downstream random forest classification with weather data
The Surface Condition Model is a two-stage approach that begins with image classification (CNN) and is followed by downstream classification model, a random forest (RF) trained on both weather data and the probability output of the CNN model. The RF model is considered “downstream” because it requires the output from the CNN, and also because it is trained on a separate dataset (training dataset 2) that is not used to train the CNN models, as shown in Fig. 3. This downstream RF model serves two purposes. First, it incorporates weather data into the modeling process, which is relevant given that the five classes are weather related. Second, by training the model on unseen observations (using training dataset 2), the RF model is able to adjust for biases in the CNN by learning when the CNN predictions are more (or less) correct and refine its prediction by considering the weather. Improvements gained from adding weather data are discussed in Section 4.1 and shown in Fig. 6.
First, the CNN model from stage 1 is used to generate predictions (inference) on training dataset 2. Probabilities are normalized to sum to one after calibrating the predicted class using isotonic regression, a common technique used to create better alignment between predicted probability and accuracy (Pedregosa et al., 2011; Jiang et al., 2011; Hollance, 2022). Next, the weather data are collected for the time and location corresponding to each image in the hand-labeled dataset (Section 2). The five probabilities and the six weather variables comprise the 11 inputs for the RF model, and the hand-labeled classification is the target variable. A random forest is selected for this stage in the modeling process, but other algorithms and hyperarameter combinations are tested, as shown in Table 4. The hyperparameters affect the complexity of each algorithm, thus affecting model performance.
Table 4: Surface Condition Model - Downstream Model Tuning. Algorithms including dense neural network (DNN), Gaussian Naive Bayes (GNB), Support Vector Machines (SVM), and Random Forest (RF), and corresponding hyperparamaters, are tested. The red text indicates the final selected model (discussed in Section 4.1).
Algorithm Hyperparameters
DNN Hidden layer nodes: [256,128,64,32,16], [128,64,32,16], [64,32,16], [32,16], [16] | Dropout: 0, 0.2, 0.4, 0.6, 0.8 | L2: 0, 0.001, 0.01, 0.1
GNB var_smoothing: 1e-12, 1e-11, 1e-10, 1e-8, 1e-7, 1e-6
Logistic max_iter: 100, 300, 500 | C: 0.01, 0.1, 1.0, 10.0, 100.0
SVM C: 1e-3, 1e-2, 0.1, 1, 10, 100 | gamma: 1e-7, 1e-5, 1e-4, 1e-3, 1e-2, 0.1, 1, 10
RF n_estimators: 25, 100, 300 | max_depth: 5, 10, 20, None | max_features: 3, 6, 10 | min_samples_leaf: 1, 5 | bootstrap: FALSE, TRUE | max_samples: 0.5, 0.75, 1
A random forest is a commonly used ML algorithm that captures nonlinear relationships in data (Auret and Aldrich, 2012). It is an ensemble of decision trees where each tree makes a sequence of decisions and ultimately places each observation into a final leaf node. The gini impurity index is used as the loss function (Géron, 2020) to train the model. The python ML package, scikit-learn (Pedregosa et al., 2011), is used to train RFs for this task. The outputs of the RF are model predicted classes and the probabilities of each class. A random forest is selected as the best performing model (Section 3.2.3), but another benefit of utilizing RFs is that feature importance can easily be implemented. Feature importance provides information about the contribution of each input variable, which can be helpful for determining which variables to include in model training, as well as for understanding which features the model relies heavily on to make its classification (see Section 4.3).
@@ 3.2.3 Model selection
Accuracy is the percentage of correct model predictions (number of correct predictions divided by the number of total observations), and recall can be thought of as a measurement of accuracy within a given labeled class (e.g., number of correct snow predictions divided by the number of observations labeled snow). Recall, precision, and F1 score are metrics that are commonly used in ML (Géron, 2020), but in this study, recall is prioritized because it is the class-specific metric that minimizes the number of misses, which is especially important for this work. As an example, it is crucial for traveler safety that the model does not miss severe snow cases, and recall is the metric that captures this goal. Additionally, since recall measures the performance relative to the number of labeled images in each class, it is an appropriate metric to use with the largely imbalanced dataset. Relying on accuracy alone is insufficient because it is heavily influenced by class imbalance. Performance of each class is important for this application, and this is accounted for by calculating the average recall across the five classes (i.e., an average of the five metrics: severe snow recall, snow recall, wet recall, dry recall, poor visibility recall). This average recall metric serves as a more balanced performance metric because each class has an equal contribution, rather than accuracy, which is heavily skewed toward the majority classes. Both metrics - accuracy and average recall - are calculated on the validation data and used for model selection by first narrowing to the top-performing models based on average recall, and then selecting the model with the highest accuracy among them.
@@ 4.1 Surface Condition Model results
For the stage 2 downstream model that incorporates weather data, five classification algorithms with varying hyperparameters are tested (see Section 3.2.2 for details). The top performing architecture for each algorithm is shown in Fig. 5B. The GNB model has the lowest performance, while the DNN, RF, and SVM algorithms all have accuracies and average recall greater than 75%. The final selected algorithm is an RF of 300 decision trees with a max depth of 10, a maximum of three features considered in each split, a minimum of five samples per leaf node, and bootstrap sampling of 0.50. This model achieves validation accuracy of 80.7%.
Fig. 5: Surface Condition Model: Validation Performance Comparison Across Models. (A) Validation accuracy of the 36 baseline CNN models are shown; this includes Architecture-Specific Top (greens), Generic Top (blues), and None (reds), and the inclusion of Augmentation (lighter shades), and No Augmentation (darker shades). (B) The top performing hyperparameter set for each of the five algorithms from the stage 2 downstream model are shown, including overall accuracy, as well as average recall metric.
@@
The final model improves predictions compared to the CNN alone, and the most significant improvements are in the snow and dry classes, which increase by 17.3 and 24.3 percentage points, respectively, as shown in Table 5. For the final model, the classes with the lowest recall are wet, poor visibility, and snow, and this is a reasonable result given that these classes border multiple classes. For example, a simplified progression of a winter weather event is: $\text{dry}\leftrightarrow\text{wet}\leftrightarrow\text{snow}\leftrightarrow\text{severe snow}$. Snow and wet appear in the middle as “transitional” phases, and therefore have more opportunity for misclassifications, as shown in Table 5. Out of all of the model’s errors, 95.5% fell into one of the adjacent classes, which shows that the model makes reasonable errors that align with difficulties that humans also encounter when labeling observations. Similarly, poor visibility is a difficult class given that there are many conditions that can lead to this class, including dense fog, heavy snow, or heavy rain.
Table 5: Class recall metrics. Performance by class for the CNN and final model are shown.
Hand-labeled class CNN Recall Final Recall Common misclassifications
Severe snow 77.2% 81.5% Snow
Snow 59.8% 77.1% Wet, severe snow
Wet 70.6% 75.9% Dry, snow
Dry 65.8% 90.1% Wet
Poor visibility 75.0% 76.8% Wet, severe snow, dry
As a comparison, the results of Khan and Ahmed (2022) show a validation accuracy of 95.2% on a sample of $300$ camera images from different sites, which is higher than the results achieved in this work. This lower accuracy could be due to a variety of factors, including 1) lower camera quality in the NYSDOT images, 2) the inclusion of more road surface classes in the NYSDOT images (e.g. severe snow and poor visibility) and 3) testing on a larger dataset of approximately 3500 held-out images. The importance of data representation is discussed in detail in Section 4.3.
@@ 4.3 Generalizability takeaways
5 Conclusions
The NYSDOT has a large network of camera images that are used to automatically classify road surface conditions. A systematic hand-labeling process known as quantitative content analysis (QCA) is followed to create a well-documented hand-labeled dataset that is used as the target variable for training ML algorithms. The Surface Condition Model predicts severe snow, snow, wet, dry, or poor visibility using CNNs and RFs trained on camera images and weather data. The Obstruction Detection Model predicts obstructed or non-obstructed using camera images and a CNN. The test accuracy of the final Surface Condition Model is 81.5%, and the Object Detection Model detects obstructed observations with a recall of 88.4%.
Labeling images from every camera site is not feasible, nor does it provide a pathway for new states or regions to implement this model. Thus, the performance of unseen camera sites is prioritized in this study. One major takeaway is that the inclusion of weather data proves to be particularly beneficial for increasing model generalizability. Additional experiments reveal that more data, specifically from new and diverse sites, improve model generalizability, especially on low-quality cameras. Thus, adding more sites through additional labeling should be considered in the future and can be done reliably given the QCA process and documentation employed in this work. Another consideration for future work is that there are other ML techniques that perform well with limited data and these should also be explored, such as co-training or discriminative self-supervised learning (Rothenberger and Diochnos, 2024).
The inclusion of low-quality sites and difficult classes like poor visibility make the task more difficult but are important because the broader purpose of this work is to provide an automated product for the NYSDOT to use operationally. Collaboration with the NYSDOT is critical for the adoption and usability of this work. The models in this study can provide a helpful tool to monitor current road surface conditions because it can update frequently (every five minutes) and vastly (at the camera level), which is of value to the NYSDOT, and can be applicable to other states and regions. Resource allocation and driver safety are crucial, and with the range of winter weather conditions that New York State experiences, this study presents a method of automated road surface classification that can offer pertinent insights during hazardous events.

### v2

@@ 3 Modeling methods
The labeled dataset and its corresponding image and weather data are used to predict road surface conditions across NYS. Classification ML algorithms, including CNNs and RFs, are trained to categorize each observation into one of five road surface classes for the Surface Condition Model. For the Obstruction Detection Model, a CNN is utilized to classify each observation as either obstructed or non-obstructed. Ensembling methods then combine predictions from multiple models trained on different datasets (Sections 2.2 and 3.4). All models use appropriately normalized and preprocessed input data. Balanced class weighting is used to address the imbalanced dataset (Table 1), which is important to encourage the model to learn all classes rather than simply learning to predict majority classes, like wet and dry.
@@ 3.1 Image classification
Images contain a wide range of colors, textures, edges, and nonlinear, complex patterns. CNNs are commonly used image models because they can automatically extract these complex features in images to learn a classification task. Spatial relationships that exist in the images are learned through filters and convolutional layers (Randy2022), and nonlinearity in the data is learned by filters and convolutional layers with activation functions.
Images are cropped to remove the top 20% of the image that often contains text (e.g., timestamps or location, examples in Table 1) or is not focused on roadways (e.g., sky or distant scenery) to balance noise removal while preserving essential road surface features within varying camera views. Images are resized to 224x224 and encoded as RGB. Cropping, resizing, and standardization of the data are preprocessing steps that are done on the fly during data loading. The model weights from the epoch achieving highest validation accuracy are saved as the product of training. Table 2 shows additional CNN modeling details.
@@ 3.2.1 CNN models and hyperparameter tuning
Six CNN architectures, which range in complexity, include DenseNet121 from densenet, InceptionV3 from inception, Mobilenet from mobilenet, ResNet50 from resnet, VGG16 from vgg16, and Xception from xception. Data augmentation is used to diversify the training set by applying transformations to the original images, including adjusting the brightness and/or contrast, and horizontally flipping the image. Augmentations are made automatically and randomly to images during model training, with replacement, to make the dataset more robust and varied (tf).
@@
Transfer learning involves using pretrained models where the backbone of the architecture (which include convolutional layers) is frozen; the weights are already learned from the commonly used ImageNet dataset (imagenet). Keras provides the architectures and pretrained weights used in this study (keras). The final layers of the architecture are the dense layers that ultimately result in the final classification, and will be referred to as the “top” of the model. The top is always trained specifically on the NYSDOT dataset. Two different architecture tops are considered: 1) the generic top (“GEN”) consists of a 2D global average pooling layer followed by three dense layers of size 1024, 512, and six (the output layer), and 2) the architecture-specific top (“AST”) consists of layers that align with the original architecture. A third option excludes transfer learning entirely, training all model weights on the NYSDOT dataset. Thus, three configurations are tested: transfer learning with GEN, transfer learning with AST, and no transfer learning.
@@
For hyperparameter tuning, L2 regularization and dropout are common regularization techniques used to reduce overfitting in CNN models (handsonml). Regularization is used on the top dense layers, and hyperparameter tuning is completed to identify the best performing L2 weight ($\lambda$) and dropout rate. The hyperparameter space (Table 3) includes four non-zero values each for dropout and L2, resulting in 16 sets. Thus, including the baseline runs where there is no regularization (dropout and L2 are 0), there are 17 hyperparameter sets tested in total.
@@
The modeling process begins with running the 36 baseline models, from which the top four are selected using the model selection criteria discussed in Section 3.2.3. The four models then undergo hyperparameter tuning, and one top performing hyperparameter set is selected for each. Thus, four final CNN models for image classification are considered and carried over into the next step of the modeling process, the downstream classification ML model, which incorporates information from the CNN as well as weather data, as discussed in section 3.2.2. The final CNN is selected following hyperparameter tuning of the downstream model.
@@
Table 3: CNN model considerations. Modeling choice sets for baseline runs and hyperparameter tuning (unitless regularization parameters) are shown. The red text indicates the final best CNN model that is selected and discussed in Section 4.1.
@@
Architecture Transf. Learn. Augment. Dropout L2 ($\lambda$)
{Densenet, Inception, Mobilenet, Resnet, VGG16, Xception} {None, AST, GEN} {Aug.,
No Aug.}
{0.0, 0.2, 0.4, 0.6, 0.8} {0.0, 1E-7, 1E-5, 0.001, 0.1}
@@ 3.2.2 Downstream random forest classification with weather data
The Surface Condition Model is a two-stage approach that begins with image classification (CNN) and is followed by downstream classification model that is trained on both weather data and the probability output of the CNN model. This stage 2 model is considered “downstream” because it requires the output from the CNN, and also because it is trained on a separate dataset (training dataset 2) that is not used to train the CNN models, as shown in Fig. 3. The downstream model serves two purposes. First, it incorporates weather data into the modeling process, which is relevant given that the five classes are weather related. Second, by training the model on unseen observations (using training dataset 2), the model is able to adjust for biases in the CNN by learning when the CNN predictions are more (or less) correct and refine its prediction by considering the weather. The improvements from adding weather data are discussed in Section 4.1 and shown in Fig. 6.
First, the CNN model from stage 1 is used to generate predictions (inference) on training dataset 2. Probabilities are calibrated using isotonic regression, a common technique used to create better alignment between predicted probability and accuracy (smooth-isotonic-regression; reliabilitycurves). Next, the weather data are collected for the time and location corresponding to each image in the hand-labeled dataset (Section 2). The five probabilities and the six weather variables comprise the 11 inputs for the downstream model, and the hand-labeled classification is the target variable. Multiple classification algorithms and hyperparameters are tested for this step of the modeling process, including dense neural network (DNN), Gaussian Naive Bayes (GNB), logistic regession, Support Vector Machines (SVM), and Random Forest (RF) (chollet_deep_learning; handsonml). Model details and results are shown in Table 4; the process of selecting models from the set of those tested is discussed in Section 3.2.3.
@@ 3.2.3 Model selection
Accuracy is the percentage of correct model predictions (number of correct predictions divided by the number of total observations), and recall can be thought of as a measurement of accuracy within a given labeled class (e.g., number of correct snow predictions divided by the number of observations labeled snow). Recall, precision, and F1 score are metrics that are commonly used in ML (handsonml), but in this study, recall is prioritized because it is the class-specific metric that minimizes the number of misses, which is especially important for this work. As an example, it is crucial for traveler safety that the model does not miss severe snow cases, and recall is the metric that captures this goal. Additionally, since recall measures the performance relative to the number of labeled images in each class, it is an appropriate metric to use with the largely imbalanced dataset. Relying on accuracy alone is insufficient because it is heavily influenced by class imbalance. Performance of each class is important for this application, and this is accounted for by calculating the average recall across the five classes (i.e., an average of the five metrics: severe snow recall, snow recall, wet recall, dry recall, poor visibility recall). This average recall metric serves as a more balanced performance metric because each class has an equal contribution, rather than accuracy, which is heavily skewed toward the majority classes. Both metrics - accuracy and average recall - are calculated on the validation data and used for model selection by first narrowing to the top-performing models based on average recall and then selecting the model with the highest accuracy among them.
@@ 4.1 Surface Condition Model results
Fig. 5: Surface Condition Model: Validation Performance Comparison Across Models. (A) Validation accuracy of the 36 baseline CNN models are shown; this includes Architecture-Specific Top (greens), Generic Top (blues), and None (reds), and the inclusion of Augmentation (lighter shades), and No Augmentation (darker shades). (B) The top performing hyperparameter set for each of the five algorithms from the stage 2 downstream model are shown, including overall accuracy, as well as the average recall metric.
Table 4: Downstream Model Tuning. Algorithms and their hyperparameters. The validation accuracy (“Acc”) of the top hyperparameter, and the standard deviation (“Std”) and a 95% confidence interval (“CI”) of accuracy across the 30 unique validation folds are shown for each algorithm. Results are also displayed visually in Figure 5. The red text indicates the final selected model.
Algo-rithm Hyperparameters Validation results
DNN Hidden layers: {[256,128,64,32,16], [128,64,32,16], [64,32,16], [32,16], [16]}, Dropout: {0,0.2,0.4,0.6,0.8}, L2: {0,0.001,0.01,0.1} Acc: 77.5% Std: 2.4% CI: (76.7%, 78.5%)
GNB var_smoothing:{1e-12, 1e-11, 1e-10, 1e-8, 1e-7, 1e-6} Acc: 69.4% Std: 3.8% CI: (68.1%, 71.9%)
Logistic max_iter:{100, 300, 500}, C:{0.01, 0.1, 1.0, 10.0, 100.0} Acc: 78.0% Std: 2.0% CI: (77.4%, 78.8%)
SVM C:{1e-3, 1e-2, 0.1, 1, 10, 100}, gamma:{1e-7, 1e-5, 1e-4, 1e-3, 1e-2, 0.1, 1, 10} Acc: 78.2% Std: 2.1% CI: (77.5%, 79.0%)
RF estimators:{25,100,300},max_depth:{5,10,20,None}, max_features:{3,6,10},min_samples_leaf:{1,5}, bootstrap:{FALSE,TRUE},max_samples:{0.5,0.75,1} Acc: 80.7% Std: 1.7% CI: (80.1%, 81.3%)
For the stage 2 downstream model that incorporates weather data, five classification algorithms with varying hyperparameters are tested, as discussed in Section 3.2.2 and shown in Table 4. The top performing architecture for each algorithm is shown in Fig. 5B. Results are also displayed in Table 4, including the standard deviation and confidence interval of accuracy across the 30 validation folds. The GNB model has the lowest skill, while the DNN, Logistic, and SVM models all have accuracies around 78%. The final selected algorithm, following model selection criteria from Section 3.2.3, is an RF of 300 decision trees with a max depth of 10, a maximum of three features considered in each split, a minimum of five samples per leaf node, and bootstrap sampling of 0.50. This model achieves validation accuracy of 80.7%.
@@
The final model improves predictions compared to the CNN alone, and the most significant improvements are in the snow and dry classes, which increase by 17.3 and 24.3 percentage points, respectively. For the final model (Table 5), the classes with the lowest recall are wet, poor visibility, and snow, and this is a reasonable result given that these classes border multiple classes. For example, a simplified progression of a winter weather event is: $\text{dry}\leftrightarrow\text{wet}\leftrightarrow\text{snow}\leftrightarrow\text{severe snow}$. Snow and wet appear in the middle as “transitional” phases, and therefore have more opportunity for misclassifications, as shown in Table 5. Out of all of the model’s errors, 95.5% fell into one of the adjacent classes, which shows that the model makes reasonable errors that align with difficulties that humans also encounter when labeling observations. Similarly, poor visibility is a difficult class given that there are many conditions that can lead to this class, including dense fog, heavy snow, or heavy rain.
@@
Table 5: Final model performance. Performance by class for the final model is shown, including recall, precision, F1 score. A one-vs-all approach is used to provide the area under the precision-recall curve for each class (PR-AUC).
Hand-labeled class Recall Precision F1-score PR-AUC Common misclassifications
Severe snow 81.5% 78.7% 80.1% .886 Snow
Snow 77.1% 68.0% 72.2% .776 Wet, severe snow
Wet 75.9% 87.8% 81.4% .914 Dry, snow
Dry 90.1% 84.4% 87.2% .943 Wet
Poor visibility 76.8% 64.2% 70.0% .698 Wet, severe snow, dry
As a comparison, the results of KHAN2022 show a validation accuracy of 95.2% on a sample of $300$ camera images from different sites, which is higher than the results achieved in this work. This lower accuracy could be due to a variety of factors, including 1) lower camera quality in the NYSDOT images, 2) the inclusion of more road surface classes in the NYSDOT images (e.g. severe snow and poor visibility) and 3) testing on a larger dataset of approximately 3500 held-out images. The importance of data representation is discussed in detail in Section 4.3.
@@ 4.3 Generalizability takeaways
5 Conclusion and future work
A large network of camera images and weather forecast data are used to predict weather-related road surface conditions with 81.5% accuracy on unseen test sites. An Obstruction Detection Model is also developed to predict obstructed image instances, correctly classifying 88.4% of obstructed cases. For operational application, the full pipeline - data preparation and model inference - runs in approximately 3.5 minutes for all camera sites statewide on a V100 GPU with 8 CPUs and 64 GB of memory, demonstrating feasibility for real-time updates every five minutes.
This study emphasizes generalizability to new and unseen sites, which is critical for operational deployment by the NYSDOT and other agencies. The inclusion of weather data proves to be particularly beneficial for increasing model generalizability through an increase of +12.1% accuracy. Including additional data from new and diverse sites further enhances generalizability, especially for low-quality cameras. The inclusion of low-quality sites and difficult classes like poor visibility make the task more difficult but are important for realistic operational deployment. The focus on generalizability, along with the robust data curation methods developed for this work (Sutterzenodo; Sutterdatasheet), provides a framework that can be extended to other agencies and locations.
Future work should include operationalization of the tool for NYSDOT through creation of a dashboard or API, including feedback from end-users to assess and improve the usefulness of the tool. Collaboration with the NYSDOT has been and will continue to be critical for the adoption and usability of this work. Expansion to other regions and states should be explored, including the addition of more labeled data. From a modeling perspective, future work should consider post-processing techniques, inclusion of additional weather variables, and testing of other modeling techniques such as recurrent neural networks, co-training or discriminative self-supervised learning (rothenberger2024meta). Overall, the models developed in this work can support frequent (every five minutes) and wide-scale monitoring of road surface conditions at the camera level, offering information valuable for resource allocation and driver safety during hazardous winter weather events.